How many mL of 0.1M HNO3 are required to neutralize exactly half of the OH ion present in 520 mL of 0.18N KOH
1
Expert's answer
2019-10-10T03:46:20-0400

n KOH =n OH = 0.52*0.18= 0.0936 mol

V HNO3 = n OH / C HNO3 = 0.0936/ 0.1= 0.936L = 936mL